# Disentangling Boltzmann brains, the time-asymmetry of memory, the H theorem, and the second law

### v2

@@ III Relation between the H theorem, Boltzmann Brains, and the second law
Another possible concern is that our argument was only semi-formal, relying on “common experience”. However, in any investigation concerning foundational issues in thermodynamics, (like whether BBs accord with our present data records), if one does not use fully formal mathematical reasoning it is extremely easy to subtly assume what one is in fact trying to prove. For example, that mistake is made in very many of the papers in the literature trying to prove the second law. Specifically, the analyses in many of these papers implicitly privilege the initial condition of a physical system (e.g., how it was set up) over its final condition (e.g., how it finishes). That asymmetry in how it specifies boundary conditions is ultimately the basis for the “derivation” of the second law. However, privileging the initial conditions in turn arises from the psychological arrow of time – which it is commonly believed to be a consequence of the asymmetry of the second law. So such a supposed derivation involves circular reasoning. The same mistake of circular reasoning has also been made in many earlier investigations of the BB hypothesis.
@@
Before presenting this formal version of our central argument, we need to say something about our terminology. Strictly speaking, the H theorem only concerns a quantity, “H”, that involves kinetic interactions among particles governed by Newtonian mechanics. So as Boltzmann stated it, that theorem does not concern other phenomena grounded in the microscopic, classical dynamics laws of the universe, like classical electromagnetism. However, the same basic reasoning Boltzmann used can be extended from the specific case of the H theorem, to concern Gibbs entropy. Accordingly, from now on we will refer to the “H theorem” when we mean the analogous result involving entropy. Taking further license with our terminology, we will use “H theorem” specifically to mean the entropy of the entire universe. Similarly, throughout this text we will often refer to “Newtonian mechanics” as shorthand for all of classical physics. With some more care about the precise meaning of “space-time events”, this shorthand could be extended even further to also include quantum mechanics and general relativity.
@@ III.1 Using probability distributions to reason about Newton’s laws
To begin, note that much of the literature (and in particular our central argument) ascribes probabilities to various possible laws of physics, e.g., to the second law. However, to meaningfully speak about a probability distribution over laws, we need to at least sketch how physical laws can be formulated in terms of sets in a sigma algebra that that probability distribution is defined over.
In this paper we use a sigma algebra $\Gamma$ defined over a state space whose elements we will call “space-time events”. To be maximally general, we don’t define that state space any further. We then define “laws” to be special kinds of partitions over the sigma algebra $\Gamma$. Since the details of how we identify elements of such partitions as laws have no bearing on our analysis, they are consigned to Appendix A.
With this issue put aside, below we investigate posterior distributions over laws given various data sets, e.g., “$P(\mbox{Newton's laws hold}\,\,|\,\,\mbox{observational data }D)$”. In this paper we will mostly be interested in such observational data sets $D$ that comprise the current state of multiple “scientific records” stored in some of our books, academic papers, computer memories, etc. Central to our analysis will be the suppositions (properly formalized via information theory) that there is high probability that every such record that we consider accurately provides the results of experiments that take place at some time other than the present.
@@
III.2 The H theorem
Now that we have elaborated how a data set providing the current state of a memory system can lead us to believe Newton’s laws, we can start to investigate the thermodynamic implications of those laws. One of the myriad contributions of Boltzmann was, of course, his argument that the H theorem follows from Newton’s laws. In modern language, that theorem concerns stochastic processes, Markov processes in particular. (See Lawler (2018); Shalizi (2007) for background on stochastic processes.) However, Boltzmann derived that theorem well before Kolmogorov formalized the notion of stochastic processes, and in particular before anyone had started to systematically investigate Markov processes. As a result, even though his mathematics deriving the H theorem was correct, his formulation of the H theorem was only semi-formal. Unfortunately, this has led to widespread failure to fully grasp the mathematics of the H theorem, and how to use it to reason formally about the second law, BBs, etc.. To undertake such reasoning one must formulate the H theorem in terms of stochastic processes.
To that end, we now present a generalized version of the H theorem, as a specific type of Markov process. While we will use the term “entropy” to mean the random variable of this process, the underlying mathematics of this kind of process is more general. Write this Markov process as ${\cal{X}}$, defined over a (compact, real-valued) state space $X$. We write $x_{t}$ for the value of that process at the time $t$, i.e., the entropy of the universe then. (Note that while $X_{t}$ the state space of a random variable, $t$ is an index, distinguishing instances of that random variable.) Refer to a distribution conditioned on $n$ time-indexed entropy value as an “$n$-time-conditioned” distribution. For example, $P(x_{3}\,|\,x_{1},x_{2})$ is a 2-time-conditioned distribution over $x_{3}$.
@@ III.2 The H theorem
$\begin{array}[]{cc}&\begin{array}[]{cc}P(x_{t+k}=a\,|\,x_{t}=b)\;=\;P(x_{t-k}=a\,|\,x_{t}=b)\\ P(x_{t+k}=a\,|\,x_{t}=b)\;=\;P(x_{t}=a\,|\,x_{t-k}=b)\\ \end{array}\end{array}$ (9)
where $a,b\in X$ and $k>0$ is arbitrary. The first equality expresses time symmetry (also called time reversal invariance) about a particular time $t$. The second expresses time translation invariance. The particular stochastic process derived in Boltzmann’s H theorem is a Boltzmann process. With $k$ implicit, we refer to the conditional distribution in Eq. 9 as the kernel that generates the Boltzmann process. (Because of the time-translation and time-symmetric invariances of the process, the arguments of the kernel are not indexed by any specific values of time.)
@@
We refer to this result as the “Boltzmann generation lemma”. Note that due to Markovanity, conditioning on an additional value $x_{t_{i}}$ has no effect on this conditional distribution if $t_{i}$ lies outside the interval $[t_{0},t_{f}]$. So the lemma provides a broad suite of $n$-time conditioned Boltzmann distributions, for $n$ arbitrarily large. [^3] Note that the relationship between Boltzmann processes and the kernel that generates them is essentially identical to the relationship between (one-dimensional) Gaussian processes and their (Gaussian) distribution kernels. In particular, like all other stochastic processes, both Boltzmann and Gaussian processes are probability distributions defined in terms of a countably infinite set of random variables (whose index is called “time” in the case of Boltzmann processes). This means that for both of them we can evaluate distributions conditioned on arbitrary finite sets of indexed values from their state space.
@@
Let ${\cal{X}}$ be a stochastic process over a state process $X$, and let :=${\cal{E}}:=\{x_{t_{1}},x_{t_{2}}),\ldots\}$ be an associated set of events. We will say that a different stochastic process ${\cal{X}}^{{\cal{E}}}$ is equivalent to the Boltzmann process ${\cal{X}}$, nailed to the (nailed) set of events ${\cal{E}}$, if
@@
for all sets of times ${\cal{T}}^{\prime}=\{t^{\prime}_{1},t^{\prime}_{2},\ldots\}$ and associated set of joint states $(x_{t^{\prime}_{1}},x_{t^{\prime}_{2}},\ldots,)$. In other words, the unconditioned distribution over joint states $(x_{t^{\prime}_{1}},x_{t^{\prime}_{2}},\ldots,)$ under the process ${\cal{X}}^{{\cal{E}}}$ is the same as that of the Boltzmann process ${\cal{X}}$ at those times conditioned on ${\cal{E}}$. Intuitively, we “nail down” some of the time-state pairs of the underlying Boltzmann process ${\cal{X}}$ to produce ${\cal{X}}^{{\cal{E}}}$. It immediately follows from the definition Eq. 11 that for all sets of events ′′:=${\cal{E}}^{\prime\prime}:=\{x_{t^{\prime\prime}_{1}},x_{t^{\prime\prime}_{2}},\ldots\}$,
$P_{{\cal{X}}^{{\cal{E}}}}(x_{t^{\prime}_{1}},x_{t^{\prime}_{2}},\ldots\,|\,{\cal{E}}^{\prime\prime})=P_{{\cal{X}}}(x_{t^{\prime}_{1}},x_{t^{\prime}_{2}},\ldots\,|\,{\cal{E}},{\cal{E}}^{\prime\prime})$ (12)
@@
In the sequel, especially when discussed nailed data sets, it will sometimes be convenient to rewrite the pair of an index $t$ and associated value $x$ as a pair $(t,x)$, or even abuse notation and writing it as $(t,x_{t})$. As shorthand, we will say that a stochastic process over a state space $X$ is singleton-equivalent if it is equivalent to a Boltzmann process nailed to a single event, $x_{t}$. As an example, the so-called “past hypothesis (PH)” is the informal idea that the second law of thermodynamics can be derived from Boltzmann’s H theorem if we condition the Boltzmann process of the universe’s entropy on the value of the universe’s entropy at the time of the Big Bang Reichenbach (1991); Wolpert (1992); Davies (1977); Rovelli (2022). Stating this more formally, the PH involves a singleton-equivalent stochastic process, where the singleton concerns the time of the Big Bang. (This description of the PH is elaborated in a fully formal manner below.)
@@
We now have the tools to precisely state what Boltzmann did show — and what he did not. Stated formally, what Boltzmann showed was that if Newtonian mechanics holds, then the H function evolves as a Markov process having a time-symmetric and time-translation invariant kernel, i.e., as a Boltzmann process. So the assumption that ${\cal{D}}_{1}$ is reliable tells us that the stochastic process governing the dynamics of H is a Boltzmann process. This clarification of what the H theorem actually says is our second contribution.
@@
Note that Boltzmann did not prove that the H function evolves as some specific singleton-equivalent stochastic process. In particular, he did not prove that it evolves as a Boltzmann process nailed to a singleton data set $S_{t_{0}}$ where $t_{0}$ is the present. Nor did he prove the PH, that the H function evolves as a Boltzmann process nailed to a singleton data set $S(t_{PH})$ where $t_{PH}$ is the Big Bang or some such. This further clarification of what the H theorem does not say is our third contribution.
@@
We need more than just a fully formal H theorem to analyze the BB hypothesis though. To properly understand that hypothesis we also need a properly motivated basis for determining just what nailed data set we should use. To do this, we can invoke the copious literature on desiderata motivating Bayesian reasoning. Many compelling arguments, ranging from DeFinetti’s Dutch book arguments Pettigrew (2020) to Cox’s axiomatization of a “calculus of reasoning under uncertainty”Van Horn (2003) to the derivation of subjective Bayesian decision theory based on Savage’s axioms Karni (2005), all establish the normative rule that whenever we perform scientific reasoning, we should only consider probability distributions that are conditioned on precisely all of the current observational data that we have, and not on anything more that that. (See Fishburn (1981) for a comprehensive though now somewhat dated review of the relevant literature on Bayesian reasoning.)
@@
III.3 The Boltzmann Brain hypothesis and the H theorem
@@
    The H theorem holds for a kernel generating a Boltzmann process ${\cal{X}}$;
@@
As pointed out above, Boltzmann proved that the first of the suppositions defining the BB hypothesis follows from Newton’s laws (and therefore from ${\cal{D}}_{1}$). However, the other two suppositions are just as crucial to have a “BB” as the term is usually considered. In particular, as mentioned above, the H theorem does not specify that the stochastic process generating the trajectory of entropy values of our universe is singleton-equivalent. So the posterior probability of the laws of the universe conditioned only on the data ${\cal{D}}_{1}$ is not peaked about the BB hypothesis, as that hypothesis is defined above. This clarification of what the H theorem does (not) imply concerning the BB hypothesis is our second contribution.
@@
In light of this clarification, is there any way we might infer the BB hypothesis from the assumption only (that ${\cal{D}}_{1}$ is true and therefore) the H theorem holds, as the standard arguments for the BB hypothesis in the literature claim to do? To answer this question, first define ${\cal{D}}_{2}$ to be some set of scientific records which have the property that if the H theorem holds (i.e., if the first supposition of the BB hypothesis holds), and ${\cal{D}}_{2}$ is reliable, then the other two suppositions of the BB hypothesis hold with high posterior probability, conditioned on both ${\cal{D}}_{1}$ and ${\cal{D}}_{2}$?
As just discussed, the H theorem by itself does not establish the second and third suppositions of the BB hypothesis. So it appears that ${\cal{D}}_{1}$ cannot play the role of ${\cal{D}}_{2}$. What if ${\cal{D}}_{2}$ were some other data, distinct from ${\cal{D}}_{1}$? In such a case the BB hypothesis would hold. We will use the term “standard argument” for the BB hypothesis to mean the assumption that there is such a ${\cal{D}}_{2}$, and so suppositions (ii) and (iii) of the BB hypothesis hold.
@@
Unfortunately, we have yet another, third data set, ${\cal{D}}_{3}$, which (via the nailed set construction rule) contradicts supposition (ii) of the BB hypothesis. It’s convenient to partition this set of data ${\cal{D}}_{3}$ into two distinct subsets, ${\cal{D}}_{3}(1)$ and ${\cal{D}}_{3}(2)$, and discuss them separately. First, whatever else we might infer concerning the entropy of the universe at various times, we certainly have at least as much data leading us to infer the value of the universe’s entropy at the present, $S_{t_{0}}$. ${\cal{D}}_{3}(1)$ is that data leading us to conclude that the entropy at the present is that particular value. (More formally, we adopt a prior such that ${\cal{D}}_{3}(1)$ is reliable, where ${\cal{D}}_{3}(1)$ implies that equality with high probability.) Having such data establishes the second supposition of the BB hypothesis. Invoking the construction rule, this means that the singleton $S_{t_{0}}$ should be added to the nailed set.
@@
The second component of ${\cal{D}}_{3}$, ${\cal{D}}_{3}(2)$, is a set of data which have the property that if the H theorem holds (and assuming a prior such that ${\cal{D}}_{3}(2)$ is reliable), that set of data would imply that at the time $t^{*}$ at or shortly after the Big Bang, the universe had some specific entropy value $S^{*}$ that is far lower than $S_{t_{0}}$. A very large number of such data sets are provided by modern observational cosmology. Invoking the construction rule, this means that the pair $(t^{*},S^{*})$ should also be added to the nailed set, in addition to $(t_{0},S_{0})$. This implies that our nailed data set should have two time-indexed entropy values in it.
@@
As we describe below, depending on the precise parameters of the associated Boltzmann process and the precise values in this two-moment nailed data set, the associated posterior expectation of the values of the universe’s entropy would not monotonically increase as we move further into our past from the present. So at least for those precise values in ${\cal{D}}_{3}$, the three suppositions that jointly lead to the BB hypothesis would not all hold. Indeed, ${\cal{D}}_{3}$ would not just contract the usual argument establishing the BB hypothesis; it would rule out the BB hypothesis itself. (See Fig. 1.)

### v3

@@ III Relation between the entropy conjecture, Boltzmann Brains, and the second law
Another possible concern is that the arguments are only semi-formal, relying on “common experience”. However, in any investigation concerning foundational issues in thermodynamics, (like whether BBs accord with our present data records), if one does not use fully formal mathematical reasoning it is extremely easy to subtly assume what one is in fact trying to prove. For example, that mistake is made in very many of the papers in the literature trying to prove the second law. Specifically, the analyses in many of these papers implicitly privilege the initial condition of a physical system (e.g., how it was set up) over its final condition (e.g., how it finishes). That asymmetry in how it specifies boundary conditions is ultimately the basis for the “derivation” of the second law. However, privileging the initial conditions in turn arises from the psychological arrow of time – which it is commonly believed to be a consequence of the asymmetry of the second law. So such a supposed derivation involves circular reasoning. The same mistake of circular reasoning has also been made in many earlier investigations of the BB hypothesis.
@@
Before presenting this formal version of our central argument, we need to say something about our terminology. Throughout this text we will often refer to “Newtonian laws” as shorthand for all of classical physics. With some more care about the precise meaning of “space-time events”, this shorthand could be extended even further to also include quantum mechanics and general relativity.
@@ III.1 Using probability distributions to reason about Newton’s laws
To begin, note that much of the literature (and in particular our central argument) ascribes probabilities to various possible laws of physics, e.g., to the second law. However, to meaningfully speak about a probability distribution over laws, we need to at least sketch how physical laws can be formulated in terms of sets in a sigma algebra that that probability distribution is defined over. In particular, to fully clarify the (in)validity of the BB hypothesis, the PH, and the second law, we need to pay scrupulous attention to how we reason from current data to probabilistic conclusions concerning the laws of the universe.
In this paper, for completeness, we provide a rigorous definition of what it means to have a probability measure over a space of possible physical laws of the universe. Our definition involves a sigma algebra $\Gamma$ defined over a state space whose elements we will call “space-time events”. For maximal generality, we don’t define that state space any further. We then define “laws” to be special kinds of partitions over the sigma algebra $\Gamma$. This is our first contribution – a fully rigorous formalization of how to define probability measures over laws of the universe. [^2] Arguably, such a formalization of probability measures over a space of laws should be central to much of modern cosmology, as well as work in the foundations of physics, Bayesian philosophy of science, and arguably even to some of epistemology. Nonetheless, it seems that no such fully rigorous formalization has been previously introduced in the literature.
However, the details of this contribution of ours, and in particular of how we identify elements of such partitions of $\Gamma$ as possible laws of the universe, do not directly arise in our analysis of the Boltzmann brain hypothesis. Accordingly, the details of our definition are consigned to Appendix A. The reader who wishes to confirm that we do not run afoul of these restrictions on the sigma algebra of $\Gamma$ are invited to consult that appendix.
With this issue resolved, we have the machinery to reason from current data to conclusions concerning the laws of the universe in a , i.e., to investigate posterior distributions over laws given various data sets, e.g., $P(\mbox{Newton's laws hold}\,|\,\mbox{observational data }D)$”. In this paper we will mostly be interested in such observational data sets $D$ that comprise the current state of multiple “scientific records” stored in some of our books, academic papers, computer memories, etc. Central to our analysis will be the suppositions (properly formalized via information theory) that there is high probability that every such record that we consider accurately provides the results of experiments that take place at some time other than the present.
@@
III.2 The entropy conjecture
Now that we have elaborated how a data set providing the current state of a memory system can lead us to believe Newton’s laws, we can start to investigate the thermodynamic implications of those laws. One of the myriad contributions of Boltzmann was, of course, his argument that the entropy conjecture follows from Newton’s laws. (More precisely of course, he argued that Newton’s laws led to the H theorem; the entropy conjecture can be viewed loosely as a generalization of that theorem.) In modern language, his argument concerns stochastic processes, Markov processes in particular. (See Lawler (2018); Shalizi (2007) for background on stochastic processes.) However, Boltzmann derived his results well before Kolmogorov formalized the notion of stochastic processes, and in particular before anyone had started to systematically investigate Markov processes. As a result, even though his mathematics behind his analysis was correct, his formulation of those results does not explicitly involve stochastic processes.
Unfortunately, this precise failure to formulate arguments concerning the dynamics of the universe’s entropy as a stochastic process has led to widespread failure to fully grasp the mathematics of the entropy conjecture, and how to use it to reason formally about the second law, BBs, etc.. To undertake such reasoning one must formulate the entropy conjecture in terms of stochastic processes.
To that end, we now present a generalized version of the entropy conjecture, as a specific class of Markov processes. While we will use the term “entropy” to mean the random variable of this process, the underlying mathematics of this kind of process is more general. Write this Markov process as $\cal{X}$}, defined over a (compact, real-valued) state space $X$. We write $x_{t}$ for the value of that process at the time $t$, i.e., the entropy of the universe then. (Note that while $X_{t}$ the state space of a random variable, $t$ is an index, distinguishing instances of that random variable.) Refer to a distribution conditioned on $n$ time-indexed entropy value as an “$n$-time-conditioned” distribution. For example, $P(x_{3}\,|\,x_{1},x_{2})$ is a 2-time-conditioned distribution over $x_{3}$.
@@ III.2 The entropy conjecture
[…]
where $a,b\in X$ and $k>0$ is arbitrary. The first equality expresses time symmetry (also called time reversal invariance) about a particular time $t$. The second expresses time translation invariance. The particular stochastic process derived in Boltzmann’s H theorem can be viewed as a kind of Boltzmann process. With $k$ implicit, we refer to the conditional distribution in Eq. 6 as the kernel that generates the Boltzmann process. (Because of the time-translation and time-symmetric invariances of the process, the arguments of the kernel are not indexed by any specific values of time.)
@@
We refer to this result as the “Boltzmann generation lemma”. Note that due to Markovanity, conditioning on an additional value $x_{t_{i}}$ has no effect on this conditional distribution if $t_{i}$ lies outside the interval $t_{0},t_{f}]$. So the lemma provides a broad suite of $n$-time conditioned Boltzmann distributions, for $n$ arbitrarily large. [^4] Note that the relationship between Boltzmann processes and the kernel that generates them is essentially identical to the relationship between (one-dimensional) Gaussian processes and their (Gaussian) distribution kernels. In particular, like all other stochastic processes, both Boltzmann and Gaussian processes are probability distributions defined in terms of a countably infinite set of random variables (whose index is called “time” in the case of Boltzmann processes). This means that for both of them we can evaluate distributions conditioned on arbitrary finite sets of indexed values from their state space.
@@
[…]
@@
[…]
@@
In the sequel, especially when discussed nailed data sets, it will sometimes be convenient to rewrite the pair of an index $t$ and associated value $x$ as a pair $(t,x)$, or even abuse notation and writing it as $(t,x_{t})$. As shorthand, we will say that a stochastic process over a state space $X$ is singleton-equivalent if it is equivalent to a Boltzmann process nailed to a single event, $x_{t}$.
As an important example, the so-called “past hypothesis (PH)” is the informal idea that the second law of thermodynamics can be derived from the entropy conjecture if we condition the Boltzmann process of the universe’s entropy on the value of the universe’s entropy at the time of the Big Bang Reichenbach (1991); Wolpert (1992); Davies (1977); Rovelli (2022). Stating this more formally, the PH involves a singleton-equivalent stochastic process, where the singleton concerns the time of the Big Bang. (This description of the PH is elaborated in a fully formal manner below.)
@@
We now have the tools to precisely state what Boltzmann did show — and what he did not. Stated formally, what Boltzmann showed was that if Newtonian mechanics holds, then the H function evolves as a Markov process having a time-symmetric and time-translation invariant kernel, i.e., as a Boltzmann process. So the assumption that ${\cal{D}}_{1}$ is reliable tells us that the stochastic process governing the dynamics of H is a Boltzmann process. This clarification of how to properly formalize the entropy conjecture is our second contribution.
@@
Note that Boltzmann did not prove that the H function evolves as some specific singleton-equivalent stochastic process. In particular, he did not prove that it evolves as a Boltzmann process nailed to a singleton data set $S_{t_{0}}$ where $t_{0}$ is the present. Nor did he prove the PH, that the H function evolves as a Boltzmann process nailed to a singleton data set $S(t_{PH})$ where $t_{PH}$ is the Big Bang or some such. This further clarification of what the entropy conjecture does not say is our third contribution.
@@
We need more than just a fully formal entropy conjecture to analyze the BB hypothesis though. To properly understand that hypothesis we also need a properly motivated basis for determining just what nailed data set we should use. To do this, we can invoke the copious literature on desiderata motivating Bayesian reasoning. Many compelling arguments, ranging from DeFinetti’s Dutch book arguments Pettigrew (2020) to Cox’s axiomatization of a “calculus of reasoning under uncertainty”Van Horn (2003) to the derivation of subjective Bayesian decision theory based on Savage’s axioms Karni (2005), all establish the normative rule that whenever we perform scientific reasoning, we should only consider probability distributions that are conditioned on precisely all of the current observational data that we have, and not on anything more that that. (See Fishburn (1981) for a comprehensive though now somewhat dated review of the relevant literature on Bayesian reasoning.)
@@
III.3 The Boltzmann Brain hypothesis and the entropy conjecture
@@
    The entropy conjecture holds for a kernel generating a Boltzmann process $\cal{X}$};
@@
As pointed out above, Boltzmann proved that the first of the suppositions defining the BB hypothesis follows from Newton’s laws (and therefore from ${\cal{D}}_{1}$). However, the other two suppositions are just as crucial to have a “BB” as the term is usually considered. For current purposes it is safe to simply assume that item (iii) of the BB hypothesis holds, by incorporating that assumption into our prior. (While cosmological arguments indicate that that assumption holds, there are subtleties that are irrelevant for current purposes.)
@@
Item (ii) of the BB hypothesis is far more problematic however. The entropy conjecture does not specify that the stochastic process generating the trajectory of entropy values of our universe is singleton-equivalent. So the posterior probability of the laws of the universe conditioned only on the data ${\cal{D}}_{1}$ is not peaked about the BB hypothesis, as that hypothesis is defined above. This clarification of what the entropy conjecture does (not) imply concerning the BB hypothesis is our fourth contribution.
@@
In light of this clarification, is there any way we might infer the BB hypothesis from the assumption only (that ${\cal{D}}_{1}$ is true and therefore) the entropy conjecture holds, as the standard arguments for the BB hypothesis in the literature claim to do? To answer this question, first define ${\cal{D}}_{2}$ to be some set of scientific records which have the property that if the entropy conjecture holds (i.e., if the first supposition of the BB hypothesis holds), and ${\cal{D}}_{2}$ is reliable, then the other two suppositions of the BB hypothesis hold with high posterior probability, conditioned on both ${\cal{D}}_{1}$ and ${\cal{D}}_{2}$. We will use the term “standard argument” for the BB hypothesis to mean the assumption that there is such a ${\cal{D}}_{2}$, and so suppositions (ii) and (iii) of the BB hypothesis hold.
Unfortunately, whether or not we have such a ${\cal{D}}_{2}$, we definitely do have yet another, third data set, ${\cal{D}}_{3}$, which (via the nailed set construction rule) contradicts supposition (ii) of the BB hypothesis. It’s convenient to partition this set of data ${\cal{D}}_{3}$ into two distinct subsets, ${\cal{D}}_{3}(1)$ and ${\cal{D}}_{3}(2)$, and discuss them separately.
@@
First, whatever else we might infer concerning the entropy of the universe at various times, we certainly have at least as much data leading us to infer the value of the universe’s entropy at the present, $S_{t_{0}}$. ${\cal{D}}_{3}(1)$ is that data leading us to conclude that the entropy at the present is that particular value. (More formally, we adopt a prior such that ${\cal{D}}_{3}(1)$ is reliable, where ${\cal{D}}_{3}(1)$ implies that equality with high probability.) Invoking the construction rule, this means that the singleton $S_{t_{0}}$ should be added to the nailed set.
@@
The second component of ${\cal{D}}_{3}$, ${\cal{D}}_{3}(2)$, is a set of data which have the property that if the entropy conjecture holds (and assuming a prior such that ${\cal{D}}_{3}(2)$ is reliable), that set of data would imply that at the time $t^{*}$ of the Big Bang, or shortly thereafter, the universe had some specific entropy value $S^{*}$ that is far lower than $S_{t_{0}}$. A very large number of such data sets are provided by modern observational cosmology. Invoking the construction rule, this means that the pair $(t^{*},S^{*})$ should also be added to the nailed set, in addition to $(t_{0},S_{0})$. This implies that our nailed data set should have two time-indexed entropy values in it.
@@
As we describe below, depending on the precise parameters of the associated Boltzmann process and the precise values in this two-moment nailed data set, the associated posterior expectation of the values of the universe’s entropy would not monotonically increase as we move further into our past from the present. So at least for those precise values in ${\cal{D}}_{3}$, the three suppositions that jointly lead to the BB hypothesis would not all hold. Indeed, ${\cal{D}}_{3}$ would not just contradict the usual argument establishing the BB hypothesis; it would in fact rule out the BB hypothesis. (See Fig. 1.)